The utility-maximizing rule is to choose the basket of goods that has the highest marginal utility of each good in the basket.


1
Expert's answer
2021-05-10T15:01:25-0400

Solution:

The correct answer is the marginal utility to price ratio equal for all goods in the basket subject to the income constraint.


The utility-maximizing rule is to choose the basket of goods that, the marginal utility to price ratio equal for all goods in the basket subject to the income constraint.

 

A consumer should spend his limited money income on the goods which provide him the most marginal utility per dollar. A consumer will be maximizing his total utility only when the ratio of "\frac{MU}{P}" is equal for all the goods. Consumers decide to allocate their money incomes so that the last dollar spent on each product purchased yields the same amount of extra marginal utility.

The algebraic statement is that consumers will allocate income in such a way that:

"\frac{MUx}{Px} = \frac{MUy}{Py}"


Where: MUx = is the marginal utility derived from good x

            Px = is the price of good x

            MUy = is the marginal utility derived from good y

            Py = is the price of good y